OpenStudy (umulas):

Hello, I need to graph y = sec(1/3x) - 5 I know that the amplitude is 1, the period is 6pi, but I don't know what else I have to do I do know that I have to replace it at y = cos(1/3x)-5 I need to find 5 parts, making 6π/4 = 3π/2 Any ideas?

OpenStudy (mathmale):

I'd suggest that you graph y=sec x first. We'll take care of the rest of your posted problem later.

OpenStudy (umulas):

It says i'm not allowed to graph the sec, so I have to do this by hand and step by step :P For part one of the x, I got x = 1, therefore y will be 0+1 -5 = (0,-4) Part two x = 3π/2, y = -1 - 5 = (3π/2, -6)

OpenStudy (umulas):

I believe I made a mistake with the quarter periods... I checked at demos making the parts being (0,-4); (3π,-6);(6π,-4) any idea on how to fix the quarter period? For some reason I keep adding 3π/2 per quarter, not 3π

OpenStudy (mathmale):

I still think you should graph y=sec x = 1 / cos x. Where the cosine is zero, the secant is undefined. Where the cosine is +1, the secant is also +1. Please do a rough graph using the Draw utility. Please clarify: what did you intend "1/3x" to signify? 1 / (3x) ? (1/3) x?

OpenStudy (umulas):